# Akshay Krishnan1, 2  Xinchen Yan1  Vincent Casser3  Abhijit Kundu1 1Google DeepMind  2Georgia Insitute of Technology  3Waymo

### v1

@@ Appendix B Additional ablations
This section provides ablations in addition to the ones provided in our paper.
Impact of pseudo labels: As explained in Section 4.5, we found that using predictions from pretrained models [60, 22] for depth and normal as pseudo-ground truth improves the depth and normal reconstruction accuracy of our joint color-depth-normal VAE. Here, we further ablate its impact on the finetuning of Orchid for image-conditioned depth and normal estimation. As shown in Table 6, we find that the monocular depth estimation accuracy of Orchid drops by 0.3 in $\delta_{1}$, while its surface normal prediction accuracy drops significantly by approximately $1\degree$ worse mean error, if the pseudo-labels were not used for finetuning.
Model Text-conditioned generation Monocular depth Monocular normals
CLIP ($\uparrow$) FID ($\downarrow$) LPIPS ($\uparrow$) AbsRel ($\downarrow$) $\delta_{1}$ ($\uparrow$) Mean ($\downarrow$) $11.25\degree$ ($\uparrow$)
Joint latent 0.316 18.74 0.764 5.7 96.9 15.2 60.6
Distentangled latents 0.312 17.11 0.769 6.4 96.0 - -
w/o pseudo labels - - - 5.7 96.6 16.12 57.9
Table 6: Ablations: We provide further metrics to ablate some key choices in our implementation. We find that: 1) using a joint latent instead of disentangled latents to encode color, depth and normals provides text-conditioned image generations with an improved CLIP-similarity and better image-conditioned depth and normal predictions on NYUv2; 2) using pseudo-labeled depth and surface normals on a large text-image dataset when finetuning Orchid for image-conditioned prediction degrades the depth and normal estimates significantly.
@@
Disentangled latent space for joint generation: Orchid uses a unified latent space for joint color-depth-normal generation. Another alternative design to enable a joint color-depth-normal latent space would be to explicitly encode all three modalities using shared/separate VAEs, and finetune the LDM to denoise a higher dimensional concatenation of all three latents. We find that while this is a feasible approach, the quality of generated images is worse than that of using a joint latent. Our hypothesis is that this is likely due to a significant mismatch is latent representations for the UNet when compared to the color image-only pretraining stage, as opposed to a joint latent space that is similar in structure to the pretrained LDM’s latent space. Quantitatively, Table 6 shows that this disentangled latents model has a lower CLIP-similarity score when evaluated on COCO captions. It does however have a lower FID and higher LPIPS, likely because it uses the same latent dimension to store color information alone. Our joint latent however is significantly better on image-conditioned prediction tasks, as it can better leverage the interdependencies between color, depth, and normals.
@@
Figure 12: Text conditioned color-depth-normal generation: We show two predictions from Orchid for each text prompt. We qualitatively compare these to the alternative: generate color, depth and normals from a separate diffusion model for each. For this baseline, we use a color-only LDM for color, and separate Marigold [25] models for depth and normals. When comparing results, please refer to our note on depth map visualization (Section C.1).
Figure 13: Text conditioned color-depth-normal generation: We show two predictions from Orchid for each text prompt. We compare the geometry predicted by our model to Marigold [25] which uses separate models for depth and normal. When comparing results, please refer to our note on depth map visualization (Section C.1).
@@
Figure 21: Joint color-depth-normal inpainting: Given color-depth-normal images with masked regions, our model inpaints them jointly. Masked-out pixels are shown with green overlays on the input images in the leftmost column. As demonstrated in the results above, inpainted outputs from Orchid look very realistic. For e.g. in top row, the inpainted pillow seems to also have a similar color theme as the other pillow and the sofa. The inpainted results are also diverse (e.g. the canoe with and without the girl in the second example).
Appendix C Qualitative results
We provide additional qualitative results and comparisons for experiments in our paper. Colormaps used to visualize the depth and surface normal predictions is shown in 10.
C.1 Note on depth map visualization
@@
C.2 Text conditioned joint generation
@@
We show color-depth-normals generated by our model for different text prompts in Figure 11. Figure 12 compares the results from our model to a baseline that uses a color-only LDM to first generate color, and then depth and normal diffusion models to generate depth and surface normals. The results from a single pass of our model are comparable to these results. Figure 13 compares the depth and normals generated by our model to those predicted by depth and normal prediction baselines on our images (generated along with depth and normals). We find that our generated depth is comparable to depth predicted by [25] while normals from our model are significantly better.
C.3 Monocular depth and normal estimation
@@
Internet images: We show more depth and normal predictions on in-the-wild images produced by Orchid in Figures 14, 15, 16, and 17. Figures 14 and 15 compare our joint predictions to those from GeoWizard [13]. We find that our depth and normals are more accurate (with fewer errors on large sections), even though GeoWizard’s predictions more detailed in many cases. In Figures 16, and 17, we compare Orchid’s predictions to Marigold [25]. We find that Orchid has better depth estimates at longer ranges, and significantly better normal estimates overall. Note that we need different Marigold weights to predict depth and normals (unlike our joint prediction model). When comparing colorized depth maps on these datasets without ground truth depth, please refer to the note in Section C.1.
@@
Zero-shot benchmark images: We show more depth and normal predictions on the zero-shot depth and normal estimation benchmarks using in Section 4 of our paper in Figures 19, 18, and 20. Figure 19 shows that Orchid is competitive with diffusion-based depth prediction baselines Marigold [25] and GeoWizard [13], while being slightly better in some cases. Both [25] and [13] have a common failure mode where depth estimates are sensitive to image discontinuities, which our model is significantly less sensitive to. Figure 18 shows that our model is significantly better are depth estimation in outdoor environments, especially at longer ranges. Figure 20 shows that our model is significantly better at surface normals estimation, particularly on objects with curved surfaces.
C.4 Joint inpainting
@@
Section 4.4 explains how our model can be used to jointly inpaint color-depth-normals. For this task, we use as input paired color, depth, and normal images, and a user-provided mask for the region to be inpainted. In cases where only a color image is available, depth and normals can be generated using the image-conditioned Orchid. We then generate the latents in the masked region, using Orchid to iteratively denoise them, while using noise-free latents encoded from the inputs for the unmasked region. This is similar to the approach proposed in RePaint [35]. We provide qualitative results in Figure 21. We show multiple inpainting results for the same input. We find that Orchid is able to jointly generate different semantically and geometrically consistent color, depth, and normals for the masked regions.
C.5 3D reconstruction from single view
@@
The image-conditioned Orchid can jointly generate depth and normals from an input image. These color, depth, and surface normals can be used to reconstruct the 3D scene using either Gaussian Splatting methods (3DGS [26], 2DGS [23]) or Poisson surface reconstruction. We provide results at: https://orchid3d.github.io.

### v2

@@ Appendix C Ablation details
This section provides details for some of the ablations provided in our paper.
@@
Unified appearance-geometry diffusion baseline with disentangled latents: Orchid uses a unified joint latent space for color-depth-normal generation. An alternative design to enable a unified color-depth-normal diffusion model would be to explicitly encode all three modalities using separate latents (all produced by the same VAE), and finetune the LDM to denoise a higher dimensional concatenation of all three latents. We find that while this is a feasible approach, the quality of generated images is significantly worse than that of using a joint latent. Our hypothesis is that this is likely due to a significant mismatch of the latent space from the color image-only pretraining stage, as opposed to a joint latent space that is similar in structure (due to the distillation loss) and dimensionality to the pretrained LDM’s latent space. Quantitatively, Table 5 in our paper shows that this disentangled latents model has a lower CLIP-similarity score when evaluated on COCO captions. It does however have a slightly higher LPIPS, likely because it uses the same latent dimension to store color information alone. Our joint latent however is significantly better on image-conditioned prediction tasks, indicating that the model is able to learn an effective joint latent representation of all three modalities.
@@
Figure 12: Text conditioned color-depth-normal generation: We show two predictions from Orchid for each text prompt. We qualitatively compare these to the alternative: generate color, depth and normals from a separate diffusion model for each. For this baseline, we use a color-only LDM for color, and separate Marigold [24] models for depth and normals. When comparing results, please refer to our note on depth map visualization (Section E.1).
Figure 13: Text conditioned color-depth-normal generation: We show two predictions from Orchid for each text prompt. We compare the geometry predicted by our model to Marigold (separate) depth and normal models [24], and to the DepthAnything-v2 + DSINE combination[64, 2]. We find Orchid’s geometry predictions to be qualitatively better, especially on structures / people in the background in the Corgi image. Color-conditional models may be inaccurate in such cases where geometry is ambiguous. When comparing results, please refer to our note on depth map visualization (Section E.1).
@@
Figure 21: Joint color-depth-normal inpainting: Given color-depth-normal images with masked regions, our model inpaints them jointly. Masked-out pixels are shown with green overlays on the input images. Inpainted outputs from Orchid look very realistic. For e.g., the edge of the wall is a continuous straight line, unlike the inpainting generated by a color-inpainting SDXL model. The inpainted results are also diverse (e.g. the table lamp, the shape of the canoe).
Figure 22: Joint color-depth-normal inpainting: (contd. from Figure 21) Our model inpaints color-depth-normals them jointly. Masked-out pixels are shown with green overlays on the input image. Inpainted outputs from Orchid are much more realistic, including geometric details such as the shape of the cradle. On the other hand, multimodal inpainting using existing baselines produce geometric artifacts and unrealistic results. When comparing results, please refer to our note on depth map visualization (Section E.1).
Appendix D Color generation quality
Method RGB generation metrics User preference (%)
CLIP ($\uparrow$) LPIPS ($\uparrow$) Aesthetics Text adherence
RGB LDM (ours-PT) 0.319 0.741 32.1 31.0
Orchid 0.316 0.764 47.4 46.9
No notable difference - - 20.5 22.1
Table 8: Quantitative evaluation and user study for RGB quality.
While the focus of our work is not to improve the quality of generated color images, we evaluated how the quality of Orchid’s text-conditioned color generations compare to the pretrained RGB-only diffusion model that we finetune from. In Table 8, we report the commonly used CLIP score and LPIPS for both models on the MS-COCO dataset, together with the findings of a user study we conducted. We generated images from both models using different captions, and asked users to pick from 3 options - Orchid’s image, the base LDM’s image, or notable difference. The users were asked to vote on two different aspects: aesthetics (overall quality of the image), text adherence (closeness to the text caption). We surveyed 40 users with 25 images each (1000 votes across both aspects in total). The quantitative metrics show that Orchid’s generations are comparable to the color-only baseline, while the user study indicates that Orchid’s generations are slightly better, with about 20% votes indicating no notable difference between the two. These metrics depend significantly on the pretraining data and color-only model being used; Orchid maintains the pretrained generation quality while enabling joint color-depth-normal generation.
D.1 Note on depth-normal redundancy
Using a joint latent for color, depth, and normals minimizes redundancy in our latent space, in comparison to using separate latents for each modality. Depth and normals are highly inter-dependent, as normals can be derived from (metric) depth. A joint latent avoids the need for separate latents, resulting in highly consistent predictions. To further validate this redundancy, we performed a PCA analysis on concatenated (separate) depth and normal latents (8 dimensions each, 1000 samples). Only 8 PCA bases (out of the full 16 dimensions) were needed to explain >> 95% variance, confirming the strong depth-normal redundancy.
Appendix E Qualitative results
We provide additional qualitative results and comparisons for the experiments in our paper. Colormaps used to visualize the depth and surface normal predictions is shown in Figure 10.
E.1 Note on depth map visualization
@@
E.2 Text conditioned joint generation
@@
We show color-depth-normals generated by our model for different text prompts in Figure 11. Figure 12 compares the results from our model to a baseline that uses a color-only LDM to first generate color, and then depth and normal diffusion models to generate depth and surface normals. The results from a single pass of our model are comparable to these results. Figure 13 compares the depth and normals generated by our model to those predicted by depth and normal prediction baselines [24, 64, 2] on our images (generated along with depth and normals). In examples regions where depth is ambiguous for generated images (e.g. background structure in Figures 12, 13), predictions from our model are qualitatively better. In other cases, our generated depth is comparable to those of baselines [24] while our normals are significantly better.
E.3 Monocular depth and normal estimation
@@
Internet images: We show more depth and normal predictions on in-the-wild images produced by Orchid in Figures 14, 15, 16, and 17. Figures 14 and 15 compare our joint predictions to those from GeoWizard [13]. We find that our depth and normals are more accurate (with fewer errors on large sections), even though GeoWizard’s predictions more detailed in many cases. In Figures 16, and 17, we compare Orchid’s predictions to Marigold [24]. We find that Orchid has better depth estimates at longer ranges, and significantly better normal estimates overall. Note that we need different Marigold weights to predict depth and normals (unlike our joint prediction model). When comparing colorized depth maps on these datasets without ground truth depth, please refer to the note in Section E.1.
@@
Zero-shot benchmark images: We show more depth and normal predictions on the zero-shot depth and normal estimation benchmarks used in Section 4 of our paper in Figures 19, 18, and 20. Figure 19 shows that Orchid is competitive with diffusion-based depth prediction baselines Marigold [24] and GeoWizard [13], while being slightly better in some cases. Both [24] and [13] have a common failure mode where depth estimates are sensitive to image discontinuities, which our model is significantly less sensitive to. Figure 18 shows that our model is significantly better are depth estimation in outdoor environments, especially at longer ranges. Figure 20 shows that our model is significantly better at surface normals estimation, particularly on objects with curved surfaces.
E.4 Joint inpainting
@@
Section 4.4 of our paper explains how our model can be used to jointly inpaint color-depth-normals. For this task, we use as input paired color, depth, and normal images, and a user-provided mask for the region to be inpainted. In cases where only a color image is available, depth and normals can be generated using the image-conditioned Orchid. We then generate the latents in the masked region, using Orchid to iteratively denoise them, while using noise-free latents encoded from the inputs for the unmasked region. This is similar to the approach proposed in RePaint [34]. We provide qualitative results in Figure 21. We show multiple inpainting results for the same input. We find that Orchid is able to generate very realistic images, with different semantically and geometrically consistent color, depth, and normals for the masked regions. We compare this to a baseline that first inpaints color (Stable-Diffusion XL-inpainting [38]), then inpaints depth (Marigold-DC [54]), and predicts normals using Marigold/Lotus [24, 18]. The baseline performs significantly worse than Orchid, with several geometric inconsistencies in the generated color image (edges of objects or walls not intersecting, mismatch in vanishing directions, etc.). It also appears more unrealistic. The baseline uses conditional prediction on the full inpainted image for normals instead of inpainting them, as there are no publicly available normal-inpainting diffusion baselines.
E.5 3D reconstruction from single view
@@
The image-conditioned Orchid can jointly generate depth and normals from an input image. These color, depth, and surface normals can be used to reconstruct the 3D scene using either Gaussian Splatting methods (3DGS [25], 2DGS [22]) or Poisson surface reconstruction. The novel-view synthesis videos of reconstructions produced from the generated color and geometry are provided on our web page https://orchid3d.github.io.
Appendix F Limitations and future work
Orchid is not without limitations. In terms of geometry prediction accuracy, there is some scope for improvement on surfaces with high frequency edges (eg.: grass, fur, or hair). Some of these undesirably smooth predictions are apparent in our qualitative results on images in-the-wild. Future work can focus on further scaling unified appearance and geometry diffusion models, incorporating more recent developments in color diffusion models such as DiTs  and flow-matching schedules. We also anticipate unified appearance-geometry diffusion models to be applied to many downstream reconstruction settings that are beyond the scope of our work: 3D scene completion, novel-view synthesis, text-conditioned full 3D generation, 3D inpainting, etc.